OpenStudy (anonymous):

ok, says to factor completely x^2+14r+49 which would be correct? (r+7)(r+7) or (r+7)^2

OpenStudy (anonymous):

(r+7)^2

OpenStudy (anonymous):

they are the same, but the second one better

OpenStudy (anonymous):

ok, thanks!

OpenStudy (anonymous):

for the same reason that when you factor \[25\] you write \[25=5^2\] rather than \[25=5\times 5\] same thing of course, but we like exponents